The potential energy (in J) of a system in one dimension is given by: U(x) =12-2x³+7x²-4x. Which are the points of stable and unstable equilibrium for this potential function? Determine the work done in moving a particle in this potential from x = 1 m to x = 2 m
1
Expert's answer
2020-02-24T11:02:43-0500

 The work done in moving a particle in this potential from x = 1 m to x = 2 m:


"W=U(2)\u2212U(1)"

"W=(12\u22122(2)^\n3\n +7(2)^\n2\n \u22124(2))\u2212(12\u22122(1)^\n3\n +7(1) ^\n2\n \u22124(1))=3 J"

For the equilibrium:


"\frac{dU}{dx}=0=14x-6x^2-4"

The points of equilibrium:


"x_1=\frac{1}{3}\ m,x_2=2\ m"

"\frac{d^2U}{dx^2}(x_1)=14-12\frac{1}{3}=10>0"

"\frac{d^2U}{dx^2}(x_2)=14-12(2)=-10<0"

x1 is the point of stable equilibrium, x2 is the point of unstable equilibrium.